### v2

@@ Abstract
However, there was no potentially satisfying theoretical explanation until Edwards and Anderson took an interest. The experimental systems involved quenched-random site-occupation disorder together with long-range separation-decaying and sign-oscillating RKKY spin-pair interactions. EA replaced this by a technically more convenient but conceptually similar model having spins on every site with Hamiltonian $H=\sum{J_{ij}{\bf{S}}_{i}.{\bf{S}}_{j}}$, where $J_{ij}$ is the coupling between spins ${\bf{S}}_{i}$ and ${\bf{S}}_{j}$, chosen randomly from a Gaussian probability distribution of zero mean and variance $J^{2}$. They then studied its statistical physics with further inspirational concepts and mathematical tools.
Statistical physics interest is in averages, over instances of the quenched parameters ($J_{ij}$), of physical observables such as the free energy $F=T\ln{Z}$, where $Z$ is the partition function, which is generally difficult, rather than $Z$ itself, which would be easier but unphysical. To deal with this, Edwards and Anderson employed a novel but irregular mathematical procedure to effect a transformation of $\ln{Z}$ into a product of $Z$s, using the formal identity ${\ln{Z}=\lim_{n\rightarrow 0}{n^{-1}\{{Z^{n}}-1}\};Z^{n}={\prod_{\alpha=1,..n}}Z_{\alpha}}$ where the $\alpha$} label ‘replicas’ (systems with the same quenched Hamiltonian but with their spin variables evolving independently of one another).

### v3

@@ Abstract
However, there was no potentially satisfying theoretical explanation until Edwards and Anderson took an interest. The experimental systems involved quenched-random site-occupation disorder together with long-range separation-decaying and sign-oscillating RKKY spin-pair interactions. EA replaced this by a technically more convenient but conceptually similar model having spins on every site (i=1Ni=1...N) with Hamiltonian $H=-\sum_{i<j}{J_{ij}{\bf{S}}_{i}.{\bf{S}}_{j}}$, where $J_{ij}$ is the coupling between nearest-neighbour spins ${\bf{S}}_{i}$ and ${\bf{S}}_{j}$, chosen randomly from a Gaussian probability distribution of zero mean and variance $J^{2}$. They then studied its statistical physics with further inspirational concepts and mathematical tools.
Statistical physics interest is in averages, over instances of the quenched parameters $J_{ij}$}, of physical observables such as the free energy $F=-T\ln{Z}$, where $Z$ is the partition function, which is generally difficult, rather than $Z$ itself, which would be easier but unphysical. To deal with this, Edwards and Anderson employed a novel but irregular mathematical procedure to effect a transformation of $\ln{Z}$ into a product of $Z$s, using the formal identity $\ln{Z}=\lim_{n\rightarrow 0}{n^{-1}\{{Z^{n}}-1}\};Z^{n}={\prod_{\alpha=1,..n}}Z_{\alpha}$} where the $\alpha$} label ‘replicas’ (systems with the same quenched Hamiltonian but with their spin variables evolving independently of one another).